# RefAgent: A Multi-agent LLM-based Framework for Automatic Software Refactoring

### v1

@@ 1. Introduction
Large-scale software systems tend to increase in complexity and become difficult to maintain as they evolve to adapt to changing requirements (LEHMAN198419). Software refactoring is a standard practice that aims to enhance code design without altering its observable behavior (Fowler1999). Neglecting refactoring is often associated with the accumulation of technical debt, leading to an increase in code smells (LACERDA2020110610; 7503738) or design flaws that violate design principles and compromise the code understandability and maintainability (YAMASHITA20132639; 1173068). These negative consequences ultimately lead to higher maintenance costs (10.1147/sj.153.0225; Chen2016).
@@
Since manual refactoring is error-prone, time-consuming, and ineffective when extensive changes are needed (FERNANDES2020106347; inproceedings0x), various semi-automated and fully-automated techniques have been proposed over the past decade (10418470; 6032586; MORALES2017236). While IDEs offer built-in refactoring options, they support only a limited set of refactorings and still require manual effort. A Microsoft survey found that 28% of developers face challenges with refactoring, especially in large codebases and ensuring correctness (6802406). Fully-automated techniques, such as search-based refactoring, offer an appealing alternative by formulating refactoring as an optimization problem to identify a refactoring sequence that enhances the program based on a defined fitness function, which can involve factors such as code smells or software quality metrics (10.1016/j.jss.2015.03.040), but they often radically change program design (articlevahid), lack behavior preservation guarantees(10.1007/s11219-019-09477-y; alomar2021preservingbehaviorsoftwarerefactoring), and support a limited range of refactoring types, and are computationally prohibitive for large projects (10.1145/2541348.2541357; Ecole2878).
@@
Recent breakthroughs in Large Language Models (LLMs) have sparked growing interest in applying generative AI to software engineering tasks such as code generation (jiang2024surveylargelanguagemodels), fault localization (10.1145/3638530.3664174), and automatic program repair (li2024hybridautomatedprogramrepair). The emergence of LLM-based commercial tools like GitHub Copilot [^1] https://github.com/features/copilot and AutoCoderRover [^2] https://autocoderover.dev/ highlights practitioners’ growing interest in LLM-powered automated software development. Developer adoption of AI is rising (daigle2024aiwave), with 76% using or planning to use it, and 92% having tried tools like Copilot. Early efforts have also explored the use of LLMs in refactoring (cordeiro2024empiricalstudycoderefactoring; liu2024empiricalstudypotentialllms), showing promise in reducing developer effort and improving design quality. However, despite this progress, fully automated refactoring remains a largely unsolved problem. Existing techniques often suffer from one or more of the following limitations: poor scalability to large codebases, insufficient coverage of refactoring types, difficulty in maintaining behavioral correctness, and brittleness when applying transformations in real-world systems (6802406; 10.1145/2541348.2541357; Ecole2878).
@@
This persistent gap highlights the need for more dynamic, context-aware, and adaptive solutions that can not only propose refactorings but also coordinate complex, multi-step workflows involving validation, correction, and iterative refinement. In this context, LLM-based agents, defined as entities that use LLMs as the cognitive core of the agent, can dynamically adapt to changing contexts, interact with tools, and execute workflows to achieve goals (he2024llmbasedmultiagentsystemssoftware), offer a promising new direction for this challenge, not just because LLMs are powerful, but because agents can be structured to simulate the reasoning and actions of skilled developers, distributing responsibilities such as code analysis, transformation, compilation, and testing across specialized roles (liu2024largelanguagemodelbasedagents). Compared to static prompt-based LLMs, agents can reason, respond to intermediate feedback (e.g., test failures or compilation errors), and adjust plans dynamically. They can generate code, reflect on its consequences, revise decisions, and collaborate with one another, mirroring the way human developers iteratively and interactively perform refactoring in practice. Furthermore, agents enable modular reasoning: by breaking the refactoring process into smaller, goal-directed tasks, which reduces cognitive and computational complexity, improving both accuracy and interpretability (wang2024agentssoftwareengineeringsurvey; xia2024agentlessdemystifyingllmbasedsoftware).
@@
Critically, multi-agent systems bring an additional advantage: coordination across specialized roles. A single LLM may hallucinate over prolonged interactions, but a system of agents, each focused on a well-defined subtask, can collaboratively handle complexity through role specialization and communication. This design aligns closely with how real-world refactoring is performed: as a sequence of discrete, dependent steps that require both autonomy and collaboration (he2024llmbasedmultiagentsystemssoftware).
@@ 2. Background and Related Work
Recent advancements in Large Language Models (LLMs) have substantially influenced various software engineering tasks (yang2024chainofthoughtneuralcodegeneration; tian2024fixinglargelanguagemodels). Particularly, in the case of software refactoring, an empirical study by Cordeiro et al. (cordeiro2024empiricalstudycoderefactoring) evaluated the refactorings generated by StarCoder2 (lozhkov2024starcoder2stackv2) against developers. They apply zero-shot and one-shot prompting and assess the unit test pass rate using the pass@k metric. Their results show that LLMs effectively reduce code smells and apply various refactoring types. While their study focused on a single LLM with a static prompt at commit-level granularity, we introduce a fully automated multi-LLM agent approach that leverages specialized LLM agents with tool-calling capabilities to retrieve context, interact with compilation and testing environments, and perform complex, context-aware refactorings without manual intervention.
Furthermore, we base our approach on the work proposed by Choi et al. (10.1007/978-3-031-64573-0_4). The authors present a single LLM approach that iteratively refactors methods identified as having high cyclomatic complexity by refining the LLM’s suggestions using the error stack trace from compilation and testing environment across 20 iterations. However, their pipeline is limited due to sequential refactoring of individual methods, while our approach refactors the projects by iterating through all classes of the project. In addition, we decentralize the process by assigning distinct roles to specialized LLM agents, each responsible for a specific subtask and environment (e.g., compilation or testing). This modular design ensures that each agent receives and processes feedback (e.g., stack traces) within its relevant execution context. Moreover, the proposed pipeline by Choi et al. lacks contextual input for the LLM to generate the refactoring. Yet, refactoring is a context-dependent problem (MORALES2017236), thus we employ a context retrieval agent to help the LLM decide and plan the refactorings effectively.
Multi-agent systems have shown promise in software maintenance by leveraging specialized agents that work collaboratively or competitively to achieve a final goal while invoking external tools to receive context and feedback to solve issues in real-world software projects. For example, agents like AutoCodeRover (zhang2024autocoderoverautonomousprogramimprovement) and Masai (arora2024masaimodulararchitecturesoftwareengineering) use static and dynamic checks to validate generated patches, while others invoke tools such as those for syntactic correctness checking, code format checking (ma2025alibabalingmaagentimprovingautomated), and vulnerability detection (nunez). However, the use of multi-agent systems for end-to-end refactoring remains largely unexplored.
@@ 3.1. Overview
In this section, we present an overview of the overall phases of RefAgent, a multi-agent framework that iteratively improves the quality of each class in a given Java project. RefAgent is designed to simulate the sequential nature of end-to-end refactoring, where the goal is to improve weakly structured code through extensive refactoring, commonly known as Root-canal refactoring (FERNANDES2020106347). The framework allows users to run the workflow autonomously to enhance the quality of a given codebase, significantly reducing the manual effort required for large-scale refactoring tasks. For a given project, a randomly-selected target class as a starting point, the workflow of RefAgent leverages specialized LLM agents for the context-aware identification of refactoring opportunities and planning as well as ensuring the compilation of the modified source code, and thorough testing to ensure that functionality is preserved for the modified target class.
Particularly, following previous work (10.1007/978-3-031-64573-0_4), RefAgent iteratively analyzes feedback across 20 iterations using external tools such as the compilation and testing environments, and integrates linguistic reflections to refine its refactoring suggestions (shinn2023reflexionlanguageagentsverbal). This iterative self-improvement mechanism enables RefAgent to dynamically adapt and enhance the quality of code transformations. The following sections provide detailed descriptions of each of our four agents.
@@ 3.2. Context-aware Planner Agent
RefAgent employs an LLM agent, the Context-aware Planner Agent, which uses a Java class as a starting point. The agent is equipped with tool-calling capabilities that allow the model to detect when one or more tools should be called and respond with the inputs that should be passed to those tools. Specifically, it can call the context-retrieval module that runs the dependency analysis tool jdeps, passing the class name and the project path as inputs, which generates a class-level dependency graph. Next, following prior work (cordeiro2024empiricalstudycoderefactoring), the agent invokes a code smell extraction tool DesigniteJava 2.5.2 (Sharma2024) that also provides software metrics detailed in Section 4.2.2. DesigniteJava then takes the path to the source Java code as input and returns a set of quality metrics (e.g, cyclomatic complexity, lack of cohesion, etc.) that the Context-aware Planner Agent will use for identifying refactoring opportunities and generating a refactoring plan for the target class. Note that within RefAgent, we do not use its code smell detection capabilities in the refactoring process itself. Instead, the planner agent leverages only the extracted code metrics to provide context for the decision-making driven by the agent’s reasoning. We ensure that the planner’s decisions about which code regions to refactor are not biased by labeled smell instances from DesigniteJava.
@@ 3.4. Compiler Agent
The Compiler Agent employs a self-reflection loop (shinn2023reflexionlanguageagentsverbal) by interacting with the Maven compiler tool to validate the syntax of the refactored code. If a compilation error occurs, the error message is sent to the LLM, which analyzes the error, generates an error summary, and forwards it to the Refactoring Generator Agent for reflection and correction. The feedback loop is set to a maximum of 20 iterations, following previous work (10.1007/978-3-031-64573-0_4) to assess the extent to which it can effectively refine the generated refactorings. The Compiler Agent’s reflection process is as follows:
@@ 3.5. Tester Agent
Code refactoring should not alter the external functionality of the code (Fowler1999). It is thus necessary to execute tests to evaluate whether the refactored code retains its original functionality. In this process, the Tester Agent uses the code search module to locate developer-written tests from the class-level dependency graph generated by jdeps and employs the automated test suite generation tool EvoSuite to generate additional regression tests, The regression tests are created based on the original program before any refactoring is applied. We use these tests in an attempt to maximize the validation of the modified source code.
@@
Finally, this workflow is autonomously executed throughout all the classes of the project. In Section 5, we evaluate the impact of RefAgent on 8 software projects using various evaluation metrics. We compare RefAgent against developers, a search-based refactoring tool named RefGen (MORALES201825), and single-agent scenarios. We also examine the performance of our framework RefAgent, using three LLMs, notably GPT-4o, StarCoder2 and DeepseekCoder.
@@ 4.1. Studied Models and Dataset
Models. We initially design RefAgent using the closed-source GPT-4o model as its core LLM. To further assess its performance, we benchmark RefAgent against two open-source models: StarCoder2-15B-instruct (lozhkov2024starcoder2stackv2) and DeepSeekCoder-33B-instruct (deepseekcoder2024). Prior work highlights StarCoder2’s effectiveness in refactoring, showing improvements in code quality and unit test pass rates using zero-shot and one-shot prompting (cordeiro2024empiricalstudycoderefactoring). The study also mitigates data leakage by filtering recent Java projects not present in the StackV2 training data (10.1145/3379597.3387487; lozhkov2024starcoder2stackv2), an approach we follow in our setup.
@@
Dataset. We randomly select eight open-source Apache Java projects, due to resource constraints, from a dataset widely used in prior research (10.1145/3379597.3387487; cordeiro2024empiricalstudycoderefactoring). We then filter out projects that are included in StarCoder2’s training dataset Stackv2 (lozhkov2024starcoder2stackv2) to remove the risk of data-leakage on that model. Unfortunately, due to the closed-source nature of GPT-4o, we cannot ensure no data leakage for that model. We therefore rely on StarCoder2 to determine the strength of our approach in a more controlled setting. The details of the selected projects, including their size and characteristics, are summarized in Table 1.
Hardware and Computational Resources. StarCoder2 experiments are conducted on a computing server with 80 GB of memory. The operating system used is Ubuntu 22.04.4 LTS. We use a GPU-accelerated setup similar to previous work (cordeiro2024empiricalstudycoderefactoring), leveraging the Nvidia A100 GPUs for efficient model inference and code generation. The GPT4o-based experiments use GPT-4o via the OpenAI API, using cloud-based inference for model execution.
@@ 4.2.1. Code Smells Extraction
Code smells are indicators of underlying design or implementation issues that may impact maintainability, readability, and overall software quality (10.5555/2755629). Since code refactoring is commonly associated with the elimination of code smells (YAMASHITA20132639), we extract code smells from the Java projects to measure the capabilities of RefAgent in reducing code smells. We use DesigniteJava 2.5.2 (Sharma2024) to extract code smells from the source code before and after the project is refactored. DesigniteJava detects 46 different types of code smells. We assess how well RefAgent mitigates code smells compared to competing approaches. we focus on the following categories:
@@ 4.2.2. Quality Metrics Computation
We use the Quality Model for Object-Oriented Design (QMOOD), proposed by the prior work (979986), which consists of a set of quality measures using the ISO 9126 specification. It defines six high-level design quality attributes: reusability, flexibility, understandability, functionality, extensibility, and effectiveness that can be calculated using 11 lower-level design metrics as detailed in Table 2. We use QMOOD to estimate the effect of the suggested refactoring solutions on quality attributes, similarly to many prior works (articlevahid; Ecole2878). Likewise, we calculate the Quality Improvement (QI) for each quality attribute using the following formula.

### v2

@@ 1. Introduction
Large-scale software systems tend to increase in complexity and become difficult to maintain as they evolve to adapt to changing requirements (Lehman, 1984). Software refactoring is a standard practice that aims to enhance code design without altering its observable behavior (Fowler et al., 1999). Neglecting refactoring is often associated with the accumulation of technical debt, leading to an increase in code smells (Lacerda et al., 2020; Yoshida et al., 2016) or design flaws that violate design principles and compromise the code understandability and maintainability (Yamashita and Counsell, 2013; van Emden and Moonen, 2002). These negative consequences ultimately lead to higher maintenance costs (Belady and Lehman, 1976; Chen et al., 2016).
@@
Since manual refactoring is error-prone, time-consuming, and ineffective when extensive changes are needed (Fernandes et al., 2020; Yang et al., 2014), various semi-automated and fully-automated techniques have been proposed over the past decade (Alharbi and Alshayeb, 2024; Fokaefs et al., 2011; Morales et al., 2017). While IDEs offer built-in refactoring options, they support only a limited set of refactorings and still require manual effort. A Microsoft survey found that 28% of developers face challenges with refactoring, especially in large codebases and ensuring correctness (Kim et al., 2014). Fully-automated techniques, such as search-based refactoring, offer an appealing alternative by formulating refactoring as an optimization problem to identify a refactoring sequence that enhances the program based on a defined fitness function, which can involve factors such as code smells or software quality metrics (Ouni et al., 2015), but they often radically change program design (Alizadeh et al., 2020), lack behavior preservation guarantees(Baqais and Alshayeb, 2020; AlOmar et al., 2021), and support a limited range of refactoring types, and are computationally prohibitive for large projects (Pinto and Kamei, 2013; Alvarado, 2017).
@@
Recent breakthroughs in Large Language Models (LLMs) have sparked growing interest in applying generative AI to software engineering tasks such as code generation (Jiang et al., 2024), fault localization (Bin Murtaza et al., 2024), and automatic program repair (Li et al., 2024). The emergence of LLM-based commercial tools like GitHub Copilot [^1] https://github.com/features/copilot and AutoCoderRover [^2] https://autocoderover.dev/ highlights practitioners’ growing interest in LLM-powered automated software development. Developer adoption of AI is rising (Daigle and Staff, 2024), with 76% using or planning to use it, and 92% having tried tools like Copilot. Early efforts have also explored the use of LLMs in refactoring (Cordeiro et al., 2024; Liu et al., 2024a), showing promise in reducing developer effort and improving design quality. However, despite this progress, fully automated refactoring remains a largely unsolved problem. Existing techniques often suffer from one or more of the following limitations: poor scalability to large codebases, insufficient coverage of refactoring types, difficulty in maintaining behavioral correctness, and brittleness when applying transformations in real-world systems (Kim et al., 2014; Pinto and Kamei, 2013; Alvarado, 2017).
@@
This persistent gap highlights the need for more dynamic, context-aware, and adaptive solutions that can not only propose refactorings but also coordinate complex, multi-step workflows involving validation, correction, and iterative refinement. In this context, LLM-based agents, defined as entities that use LLMs as the cognitive core of the agent, can dynamically adapt to changing contexts, interact with tools, and execute workflows to achieve goals (He et al., 2024), offer a promising new direction for this challenge, not just because LLMs are powerful, but because agents can be structured to simulate the reasoning and actions of skilled developers, distributing responsibilities such as code analysis, transformation, compilation, and testing across specialized roles (Liu et al., 2024b). Compared to static prompt-based LLMs, agents can reason, respond to intermediate feedback (e.g., test failures or compilation errors), and adjust plans dynamically. They can generate code, reflect on its consequences, revise decisions, and collaborate with one another, mirroring the way human developers iteratively and interactively perform refactoring in practice. Furthermore, agents enable modular reasoning: by breaking the refactoring process into smaller, goal-directed tasks, which reduces cognitive and computational complexity, improving both accuracy and interpretability (Wang et al., 2024; Xia et al., 2024).
@@
Critically, multi-agent systems bring an additional advantage: coordination across specialized roles. A single LLM may hallucinate over prolonged interactions, but a system of agents, each focused on a well-defined subtask, can collaboratively handle complexity through role specialization and communication. This design aligns closely with how real-world refactoring is performed: as a sequence of discrete, dependent steps that require both autonomy and collaboration (He et al., 2024).
@@ 2. Background and Related Work
Recent advancements in Large Language Models (LLMs) have substantially influenced various software engineering tasks (Yang et al., 2024; Tian et al., 2024). Particularly, in the case of software refactoring, an empirical study by Cordeiro et al. (Cordeiro et al., 2024) evaluated the refactorings generated by StarCoder2 (Lozhkov et al., 2024) against developers. They apply zero-shot and one-shot prompting and assess the unit test pass rate using the pass@k metric. Their results show that LLMs effectively reduce code smells and apply various refactoring types. While their study focused on a single LLM with a static prompt at commit-level granularity, we introduce a fully automated multi-LLM agent approach that leverages specialized LLM agents with tool-calling capabilities to retrieve context, interact with compilation and testing environments, and perform complex, context-aware refactorings without manual intervention.
Furthermore, we base our approach on the work proposed by Choi et al. (Choi et al., 2024). The authors present a single LLM approach that iteratively refactors methods identified as having high cyclomatic complexity by refining the LLM’s suggestions using the error stack trace from compilation and testing environment across 20 iterations. However, their pipeline is limited due to sequential refactoring of individual methods, while our approach refactors the projects by iterating through all classes of the project. In addition, we decentralize the process by assigning distinct roles to specialized LLM agents, each responsible for a specific subtask and environment (e.g., compilation or testing). This modular design ensures that each agent receives and processes feedback (e.g., stack traces) within its relevant execution context. Moreover, the proposed pipeline by Choi et al. lacks contextual input for the LLM to generate the refactoring. Yet, refactoring is a context-dependent problem (Morales et al., 2017), thus we employ a context retrieval agent to help the LLM decide and plan the refactorings effectively.
Multi-agent systems have shown promise in software maintenance by leveraging specialized agents that work collaboratively or competitively to achieve a final goal while invoking external tools to receive context and feedback to solve issues in real-world software projects. For example, agents like AutoCodeRover (Zhang et al., 2024) and Masai (Arora et al., 2024) use static and dynamic checks to validate generated patches, while others invoke tools such as those for syntactic correctness checking, code format checking (Ma et al., 2025), and vulnerability detection (Nunez et al., 2024b). However, the use of multi-agent systems for end-to-end refactoring remains largely unexplored.
@@ 3.1. Overview
In this section, we present an overview of the overall phases of RefAgent, a multi-agent framework that iteratively improves the quality of each class in a given Java project. RefAgent is designed to simulate the sequential nature of end-to-end refactoring, where the goal is to improve weakly structured code through extensive refactoring, commonly known as Root-canal refactoring (Fernandes et al., 2020). The framework allows users to run the workflow autonomously to enhance the quality of a given codebase, significantly reducing the manual effort required for large-scale refactoring tasks. For a given project, a randomly-selected target class as a starting point, the workflow of RefAgent leverages specialized LLM agents for the context-aware identification of refactoring opportunities and planning as well as ensuring the compilation of the modified source code, and thorough testing to ensure that functionality is preserved for the modified target class.
Particularly, following previous work (Choi et al., 2024), RefAgent iteratively analyzes feedback across 20 iterations using external tools such as the compilation and testing environments, and integrates linguistic reflections to refine its refactoring suggestions (Shinn et al., 2023). This iterative self-improvement mechanism enables RefAgent to dynamically adapt and enhance the quality of code transformations. The following sections provide detailed descriptions of each of our four agents.
@@ 3.2. Context-aware Planner Agent
RefAgent employs an LLM agent, the Context-aware Planner Agent, which uses a Java class as a starting point. The agent is equipped with tool-calling capabilities that allow the model to detect when one or more tools should be called and respond with the inputs that should be passed to those tools. Specifically, it can call the context-retrieval module that runs the dependency analysis tool jdeps, passing the class name and the project path as inputs, which generates a class-level dependency graph. Next, following prior work (Cordeiro et al., 2024), the agent invokes a code smell extraction tool DesigniteJava 2.5.2 (Sharma, 2024) that also provides software metrics detailed in Section 4.2.2. DesigniteJava then takes the path to the source Java code as input and returns a set of quality metrics (e.g, cyclomatic complexity, lack of cohesion, etc.) that the Context-aware Planner Agent will use for identifying refactoring opportunities and generating a refactoring plan for the target class. Note that within RefAgent, we do not use its code smell detection capabilities in the refactoring process itself. Instead, the planner agent leverages only the extracted code metrics to provide context for the decision-making driven by the agent’s reasoning. We ensure that the planner’s decisions about which code regions to refactor are not biased by labeled smell instances from DesigniteJava.
@@ 3.4. Compiler Agent
The Compiler Agent employs a self-reflection loop (Shinn et al., 2023) by interacting with the Maven compiler tool to validate the syntax of the refactored code. If a compilation error occurs, the error message is sent to the LLM, which analyzes the error, generates an error summary, and forwards it to the Refactoring Generator Agent for reflection and correction. The feedback loop is set to a maximum of 20 iterations, following previous work (Choi et al., 2024) to assess the extent to which it can effectively refine the generated refactorings. The Compiler Agent’s reflection process is as follows:
@@ 3.5. Tester Agent
Code refactoring should not alter the external functionality of the code (Fowler et al., 1999). It is thus necessary to execute tests to evaluate whether the refactored code retains its original functionality. In this process, the Tester Agent uses the code search module to locate developer-written tests from the class-level dependency graph generated by jdeps and employs the automated test suite generation tool EvoSuite to generate additional regression tests, The regression tests are created based on the original program before any refactoring is applied. We use these tests in an attempt to maximize the validation of the modified source code.
@@
Finally, this workflow is autonomously executed throughout all the classes of the project. In Section 5, we evaluate the impact of RefAgent on 8 software projects using various evaluation metrics. We compare RefAgent against developers, a search-based refactoring tool named RefGen (Morales et al., 2018), and single-agent scenarios. We also examine the performance of our framework RefAgent, using three LLMs, notably GPT-4o, StarCoder2 and DeepseekCoder.
@@ 4.1. Studied Models and Dataset
Models. We initially design RefAgent using the closed-source GPT-4o model as its core LLM. To further assess its performance, we benchmark RefAgent against two open-source models: StarCoder2-15B-instruct (Lozhkov et al., 2024) and DeepSeekCoder-33B-instruct (DeepSeek, 2024). Prior work highlights StarCoder2’s effectiveness in refactoring, showing improvements in code quality and unit test pass rates using zero-shot and one-shot prompting (Cordeiro et al., 2024). The study also mitigates data leakage by filtering recent Java projects not present in the StackV2 training data (Claes and Mäntylä, 2020; Lozhkov et al., 2024), an approach we follow in our setup.
@@
Dataset. We randomly select eight open-source Apache Java projects, due to resource constraints, from a dataset widely used in prior research (Claes and Mäntylä, 2020; Cordeiro et al., 2024). We then filter out projects that are included in StarCoder2’s training dataset Stackv2 (Lozhkov et al., 2024) to remove the risk of data-leakage on that model. Unfortunately, due to the closed-source nature of GPT-4o, we cannot ensure no data leakage for that model. We therefore rely on StarCoder2 to determine the strength of our approach in a more controlled setting. The details of the selected projects, including their size and characteristics, are summarized in Table 1.
Hardware and Computational Resources. StarCoder2 experiments are conducted on a computing server with 80 GB of memory. The operating system used is Ubuntu 22.04.4 LTS. We use a GPU-accelerated setup similar to previous work (Cordeiro et al., 2024), leveraging the Nvidia A100 GPUs for efficient model inference and code generation. The GPT4o-based experiments use GPT-4o via the OpenAI API, using cloud-based inference for model execution.
@@ 4.2.1. Code Smells Extraction
Code smells are indicators of underlying design or implementation issues that may impact maintainability, readability, and overall software quality (Suryanarayana et al., 2014). Since code refactoring is commonly associated with the elimination of code smells (Yamashita and Counsell, 2013), we extract code smells from the Java projects to measure the capabilities of RefAgent in reducing code smells. We use DesigniteJava 2.5.2 (Sharma, 2024) to extract code smells from the source code before and after the project is refactored. DesigniteJava detects 46 different types of code smells. We assess how well RefAgent mitigates code smells compared to competing approaches. we focus on the following categories:
@@ 4.2.2. Quality Metrics Computation
We use the Quality Model for Object-Oriented Design (QMOOD), proposed by the prior work (Bansiya and Davis, 2002), which consists of a set of quality measures using the ISO 9126 specification. It defines six high-level design quality attributes: reusability, flexibility, understandability, functionality, extensibility, and effectiveness that can be calculated using 11 lower-level design metrics as detailed in Table 2. We use QMOOD to estimate the effect of the suggested refactoring solutions on quality attributes, similarly to many prior works (Alizadeh et al., 2020; Alvarado, 2017). Likewise, we calculate the Quality Improvement (QI) for each quality attribute using the following formula.